OpenStudy (briellshaw):

Simplify 67.13 - 2.32 * (5.09 - 3.09)^4 30.01 1,036.96 46.01 48.57

OpenStudy (unklerhaukus):

\[67.13 - 2.32 \times (5.09 - 3.09)^4\] simplify the terms in the brackets first

OpenStudy (unklerhaukus):

\(5.09-3.09 = . . .\)

OpenStudy (briellshaw):

2^4=16

OpenStudy (unklerhaukus):

so you have \[67.13-2.32\times16\]

OpenStudy (briellshaw):

30.01

OpenStudy (unklerhaukus):

how did you do that multiplication?

OpenStudy (briellshaw):

my dad gave me the answer

OpenStudy (unklerhaukus):

\[2.32\times 16 = 2\times16+0.1\times3\times16+0.01\times2\times16\\\qquad\qquad=32+4.8+0.32=37.12\] \[67.13-2.32\times16=67.13-37.12=30.01\]

OpenStudy (unklerhaukus):

How did your dad do the multiplication?

OpenStudy (briellshaw):

idk